the average hour spent using the computer by 20 senior high school students during online class is 7 hour with a standard deviation of 2 hours. construct a 98% confidence interval of average hour spent by all senior high school students. identify the width of the interval

1
Expert's answer
2022-02-22T00:49:29-0500

The critical value for "\alpha = 0.02" and "df = n-1 = 19" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.539483."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times \dfrac{s}{\sqrt{n}},\bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(7-2.539483\times\dfrac{2}{\sqrt{20}},7+2.539483\times\dfrac{2}{\sqrt{20}})"

"=(5.8643, 8.1357)"

Therefore, based on the data provided, the 98% confidence interval for the population mean is "5.8643<\mu<8.1357," which indicates that we are 98% confident that the true population mean "\mu" is contained by the interval "(5.8643, 8.1357)."

The width of the interval is


"W=2(t_c\times \dfrac{s}{\sqrt{n}})=2.2714"